has 5 chocolate chip, 2 sugar, and 3 peanut butter cookies. you randomly grab 3. let x be the number of chocolate cookies that you have. find the probability distribution, mean, and standard
Answer #1

Probability of a chocolate cookie= 5/10=0.5= P (success)=p

Let x be the number of chocolate cookies you have

X~ Binomial(n=3, p=0.5)

PROBABILITY DISTRIBUTION

X 0 1 2 3
P(X=x)

3C0(1-0.5)3

= 0.125

3C1(0.5)(1-0.5)2

0.375

3C2(0.5)2(1-0.5)

0.375

3C3(0.5)3

0.125

MEAN= np= 3* 0.5= 1.5

Standard dev= =​​​​​​​ = 0.866